OpenStudy (anonymous):

Find the value of $1,300 invested at 4.2% interest compounded continuously for 5 years, 9 months. Round your final answer to the nearest hundredth, or cent. Show your work. Use e= 2.718.

OpenStudy (anonymous):

so far I have 1300* e^(0.042)(3/4)

OpenStudy (anonymous):

im just wondering if that equation is correct

OpenStudy (anonymous):

because 3/4 of an yr. is 9 mos.

OpenStudy (anonymous):

hello?

ganeshie8 (ganeshie8):

hello :) correct, also dont forget to include 5 years

ganeshie8 (ganeshie8):

5 years 9 months 5 years + 9 months 5 years + 3/4 year 5 + 3/4 year 23/4 years

OpenStudy (anonymous):

yea sorry that's what I meant :)

OpenStudy (anonymous):

i got $1655.10 for an answer?

ganeshie8 (ganeshie8):

looks correct, good work :)

OpenStudy (anonymous):

ok & what abt this one?

OpenStudy (anonymous):

Use the Properties of Logarithms to solve the equation for x. Round your answer to the nearest hundredth. Show your work. e^(5x)+2=6

ganeshie8 (ganeshie8):

\(\large e^{(5x)} +2=6\)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

subtract 2 first?

ganeshie8 (ganeshie8):

thats good move :) keep going

OpenStudy (anonymous):

e^(5x)=4

OpenStudy (anonymous):

then u do.. log(4) divided by

ganeshie8 (ganeshie8):

yes, take ln both sides

OpenStudy (anonymous):

ln(4)/ln(e^5 )

ganeshie8 (ganeshie8):

\(\large e^{(5x)}=4 \) take ln both sides \(\large \ln e^{(5x)}= \ln 4 \) \(\large 5x \ln e= \ln 4 \) \(\large 5x = \ln 4 \) \(\large x = \frac{\ln 4}{5}\)

ganeshie8 (ganeshie8):

plugit in your calculator

ganeshie8 (ganeshie8):

what you got is also correct :) good work !!

ganeshie8 (ganeshie8):

i gotta run, cya

OpenStudy (anonymous):

i got .278

OpenStudy (anonymous):

ok bye, thanks

ganeshie8 (ganeshie8):

i got .277

ganeshie8 (ganeshie8):

rounding to the nearest hundreth it wud be :- .28

OpenStudy (anonymous):

yup. thanks :)